if the lengths of a pair of opposite sides of rectangle were doubled and the other pair of opposite sides were each decreased by 14 meters,the new rectangle would be a square with a perimeter of 84 meters. find the rectangle's original dimension..
Answer by AAfter Search(61)

Let the original length and breadth of the rectangle be l m and b m respectively.
On changing, the new dimensions become l - 14 and 2b.
Given, l - 14 = 2b = 84/4 = 21
=> l - 14 = 21
=> l = 21 + 14 = 35
2b = 21 m
b = 21/2 = 10.5m
Hence, the original length and breadth of the rectangle are 35 m and 10.5 m